# spacing=nonfrench

### v2

@@ II.2.4 Optimal conditions
Overall, this implies that each process should be fast and symmetrical at mean temperature $T_{0}$. The corresponding simultaneous optimization of $\alpha$ and $\rho$ implies maximization of the dissipated entropy $\sigma$, and thus of the energy extracted from the temperature gradient.
Optimization can also be based on the efficiency of the global system in extracting and using energy from the temperature gradient, as assessed by the entropic yield (Eqs. (25aibc)). Most of the entropy is dissipated for the transport of matter between the compartments for $\lambda^{\prime}\delta>\rho$, and for the nonequilibrium reaction within each compartment for $\lambda^{\prime}\delta<\rho$.
Finally, it is possible to directly optimize a specific process, regardless of the intensity of other processes. The optimization of the reaction process via $\sigma_{r}$ (see Eqs. (25aibf) and Fig. S1 in the SI) implies that:
@@
A similar reasoning can be used to optimize the transport process via $\sigma_{d}$ (see Eqs. (25aibg)):
@@ II.3.1 Model description
The spatial extension of the system was considered by introducing the chemical diffusion of all compounds. We limited our study to closed one-dimensional systems. A stationary temperature gradient $\theta(\bar{x})$ was imposed over the entire system. This corresponds to an unstirred system, with free internal diffusion of chemical compounds, without matter exchange with the surroundings. The system was described using a set of partial differential equations (see Eqs. (25aibh)), which was numerically solved for a wide range of parameters, with a focus on determining its steady state (see Appendix A). The characteristic parameter $l_{0}$ must be introduced to account for the spatial extension, describing a one dimensional system:
$\bar{x}=\frac{x}{l_{0}}\in[0,1]$ (14)
Chemical diffusion is described based on the diffusion parameter $d_{i}$, which is nondimensionalized by the characteristic parameter $D_{0}$ corresponding to the average diffusion constant of the chemical reactants. Consistently with the description of the chemical exchange kinetics of the previous model, chemical diffusion is described as a function of the geometric mean of the diffusion constant $d$ and the exchange ratio $\lambda$ (see Eqs. (25ag)). The characteristic time $t_{0}$ is additionally fixed by $l_{0}$ and $D_{0}$ (see Table. 1). All parameters corresponding to chemical reactions are defined as in the stirred compartment model; however, they are now functions of position $\bar{x}$.
@@
$w$ $=\sqrt{\frac{\lambda^{\prime}d}{2\rho}}$ (15)
@@
  • The Soret effect was neglected, which is expected to be satisfactory for systems based on small molecular reactants[37].
@@
When $w>0.5$, the reaction-diffusion length is larger than the sizes of the cold and warm regions $\bar{x}\in[0,0.5]$ and $\bar{x}\in[0.5,1]$. Both are homogeneous; the full system is then equivalent to the stirred-compartment model. This system is characterized by fast diffusion throughout the system, whereas the chemical reaction flux and dissipation are evenly distributed throughout the entire system (see Fig. 3B).
When $w<0.5$, an exponential decrease is observed in the values of $\varphi$ and $\alpha$, from their boundary values at $\bar{x}=0.5$ to $x=0$ or $x=1$, with a half-length equal to $w$. If $w\ll 1$, the activity of the system is restricted to a short region close to the boundary, most of the system being in equilibrium and characterized by $\varphi\approx 0$, $\alpha\approx 0$, $\sigma\approx 0$ (see Fig. 3D).
In both cases, a reaction-diffusion cycle is established, with chemical reaction and diffusion occurring simultaneously in the same spatial region, at the system boundary (see Fig. 3A-C).
The spatial extension of chemical activity in each region becomes asymmetric when the intensity of the temperature gradient is sufficiently high. This can be described by considering the difference in the kinetic rate in each region based on Eq. (25aea)-(25aeb), as:
$\rho_{\textrm{cold}}$ $=\rho e^{-\beta\frac{\theta_{g}}{1-\theta_{g}}}$ $\text{for}~{}\bar{x}\in[0,0.5]$ (16)
$\rho_{\textrm{warm}}$ $=\rho e^{\beta\frac{\theta_{g}}{1+\theta_{g}}}$ $\text{for}~{}\bar{x}\in[0.5,1]$ (17)
@@
When $\beta\theta_{g}>1$, the system becomes highly nonlinear, with a large behavioral asymmetry between each region. Typically, situations with $w_{\text{cold}}>1$ and $w_{\text{warm}}\ll 1$ can be obtained, with a homogeneous system in the cold zone and restricted to a short region in the warm zone (see Fig. 3E).
@@
  • Fast-diffusion systems (i.e. $w\gg 1$, see Fig. 3G) are characterized by homogeneous concentration profiles. This induces a linear variation in $\alpha$ and $\varphi$, reproducing the shape of the temperature profile, and thus a parabolic shape for $\sigma=\alpha\varphi$. Consequently, the system is more active at the extremities of the system, with $\alpha=0$, $\varphi=0$ and $\sigma=0$ close to $\bar{x}=\nicefrac{{1}}{{2}}$.
  • In fast reaction systems (i.e. $w\ll 1$, see Fig. 3I), the central inactive region characterized by $\alpha\approx 0$, $\varphi\approx 0$ and $\sigma\approx 0$ becomes much larger. The chemical reaction flux is restricted at each extremity of the system, in a region of length $w_{\text{cold}}$ at $\bar{x}\approx 0$ and $w_{\text{warm}}$ at $\bar{x}\approx 1$. In turn, chemical transport spreads, linking these two extremities.
This general behavior implies a spatial separation de facto of the different chemical processes. Three regions can be identified: cold and warm regions at each extremity, of half-lengths equal to $w_{\text{cold}}$ and $w_{\text{warm}}$, separated by a central region of length approximately equal to $1-2w$. The cold and warm regions are dominated by chemical reactions ($\sigma\approx\sigma_{r}$), and the central region is dominated by chemical transport ($\sigma\approx\sigma_{d}$). Thus, a spatially extended reaction-diffusion cycle is set in motion; the compounds are transported across the entire system between each chemically active region, the central transport region shrinking with $w$ (see Fig. 3F). This is in sharp contrast with the Heaviside gradient, where both chemical reactions and transport are effective at the same spatial locations, close to the temperature boundary (see Fig. 3A).
A direct consequence of this chemical transport between spatially separated regions is a sharp loss of efficiency of the dissipative processes for small $w$: the boundary values of $\alpha$ and $\varphi$ are no longer scale independent as observed with Heaviside profiles, but decrease proportionally to $w$ when $w\ll 1$. Thus, slow-diffusion systems (or large systems) are much closer to equilibrium for linear temperature gradients than for Heaviside gradients.
Finally, the high asymmetrization of the profile induced by high $\beta\theta_{g}$ values leads to nontrivial and complex flux profiles (see Fig. 3J). The central region around $\bar{x}=\nicefrac{{1}}{{2}}$ is characterized by a non-negligible positive flux, corresponding to a low-activity extension of the warm region. Consequently, the spatial position of the zero flux is localized at $\bar{x}<\nicefrac{{1}}{{2}}$; the cold region of the negative flux shrinks and the warm region of the positive flux stretches.
@@
Figure 4: Mixed temperature profile. A precise localization of the reaction diffusion cycle is achieved by a tailored temperature gradient profile. Diffusion and reaction processes are physically separated. A: temperature, with constant temperature $-\theta_{g}$ until ①, linear variation of $\theta$ from ① to ②, and constant temperature $+\theta_{g}$ from ②; B, G, diffusion flux $j_{2}$; C, F, normalized reaction flux $\varphi/\varphi_{0}$, with $\varphi_{0}=2\beta\gamma\theta_{g}\rho$; D, H, diffusion flux $j_{1}$; E, concentration $u_{2}$. Model parameters: $\kappa=1$, $\lambda=1$, $\gamma=0.2$, $d=1$, $\beta=1$, $\theta_{g}=10^{-1}$, A-D: $\rho=10^{3}$, E-H: $\rho\in[10^{-1},10^{3}]$.
The behaviors of both the linear and Heaviside profiles can be recognized (see Fig. 4). The reaction-diffusion cycles are now in motion, with two chemical reaction zones located at $\bar{x}_{1}$ and $\bar{x}_{2}$, connected by an active chemical transport region across $\bar{x}\in[\bar{x}_{1},\bar{x}_{2}]$.
It appears that controlling the precise shape of an imposed temperature gradient enables precise control of the shape of the induced reaction-diffusion cycle. Steady dissipative chemical reactions can be localized in regions characterized by temperature inflection points (i.e., high values of $\nicefrac{{\partial^{2}T}}{{\partial x^{2}}}$), whereas chemical transport can be sustained between these active regions along temperature gradients (i.e., following high values of $\nicefrac{{\partial T}}{{\partial x}}$.
II.3.3 General description
@@
The values of the steady chemical flux $\varphi$, chemical force $\alpha$ and entropy dissipation $\sigma$ can be expressed as functions of the following characteristic values:
$\alpha_{0}$ $=2\beta\gamma\theta_{g}$ (18a)
$\varphi_{0}$ $=\kappa^{\prime}\beta\gamma\theta_{g}\rho$ (18b)
$\sigma_{0}$ $=2\kappa^{\prime}\beta^{2}\gamma^{2}\theta_{g}^{2}\rho$ (18c)
@@
Table 2: First order values for $\varphi$, $\alpha$ and $\sigma$ for low values of $\beta\theta$, as identified from numerical models (see Fig. S3 in SI). The max index correspond to maximal values located at the cold/warm interface at $\bar{x}=\nicefrac{{1}}{{2}}$ for the Heaviside profile, at the system extremities $\bar{x}=0$ and $\bar{x}=1$ for the linear profile, and the inflection point of the temperature profile at $\bar{x}=\nicefrac{{1}}{{4}}$ and $\bar{x}=\nicefrac{{3}}{{4}}$. The integral values correspond to the integration of the corresponding variable on the full system $\bar{x}\in[0,1]$ for the $\sigma$ variables, or on the region restricted to the positive values of $\varphi$ and $\alpha$ (i.e. $\bar{x}\in[\bar{x}_{0},1]$, where $\varphi>0$ or $\alpha>0$ for $x>\bar{x}_{0}$).
@@
$\alpha_{0}$ $=\frac{2}{\kappa^{\prime}\rho}\varphi_{0}$ (19)
$\sigma_{0}$ $=\alpha_{0}\varphi_{0}$ (20)
In the limit cases $w\ll 1$ and $w\gg 1$, and when the system is close to equilibrium (i.e. $\beta\theta_{g}<1$), the ratio of each variable to its reference value is only dependent on the reaction-diffusion length $w$ as defined in Eq. (15). The corresponding values are summarized in Table 2 (see Fig. S3 in SI for a graphical representation).
The situation $w\gg 1$ (that is, small systems) is essentially independent of scale. This implies that, as long as the dimension of the system is smaller than the length of the reaction-diffusion, the steady-state behavior is essentially described by the characteristic values of Eqs. (18), modulated by a constant factor that depends on the shape of the temperature gradient profile. Diffusion dissipation is the only varying parameter; it is proportional to $w^{-2}$, and thus is essentially negligible for small systems. This situation is the optimal case for extracting energy from the temperature gradient (by maximizing $\sigma_{t}$, as well as driving nonequilibrium chemical reaction (c=by maximizing both $\sigma_{r}$ and $y_{r}$).
@@
The intermediate case $w\sim 1$ is characterized by the maximal value of $\sigma_{r}$, that is, by optimal nonequilibrium transport. In this regime, the chemical reaction is also important, with a dissipation of entropy equally shared in both processes for $w=1$. This situation is thus optimal for observing nonequilibrium cycles of reaction-diffusion, as both diffusion and reaction can be simultaneously efficient.
@@ II.3.4 System scaling
$\mathcal{L}_{0}$ $=\sqrt{\frac{D_{1}D_{2}}{k_{+,0}D_{2}+k_{-,0}D_{1}}}$ (21)
@@
$\mathcal{L}_{0}$ $=\sqrt{\frac{D_{0}}{2k_{0}}}$ (22)
The system spatial dimension $l_{0}$ should be adjusted accordingly. Optimal energy extraction and nonequilibrium chemical reactions are obtained for $l_{0}<\mathcal{L}_{0}$, optimal chemical transport over large distances for $l_{0}>\mathcal{L}_{0}$, and the optimal reaction-diffusion cycle for $l_{0}\sim\mathcal{L}_{0}$,
@@
Reversible peptide exchange reaction mediated by N-methylcysteine is a good candidate for being driven by temperature gradients[38]. It possesses high activation energies, and its equilibrium constant is $K=1$ close to the ambient temperature[39]:
Its thermokinetic characteristics are:
$E_{a}^{+}$ $=55\times 10^{3}~{}\text{J}.\text{mol}^{-1}$ (24a)
$E_{a}^{-}$ $=43\times 10^{3}~{}\text{J}.\text{mol}^{-1}$ (24b)
$\Delta_{r}H_{0}$ $=12\times 10^{3}~{}\text{J}.\text{mol}^{-1}$ (24c)
$k_{0}^{+}=k_{0}^{-}$ $=1.4\times 10^{-3}~{}\text{M}^{-1}.\text{s}^{-1}$ (24d)
$\text{at}\qquad T_{0}$ $=314~{}\text{K}$ (24e)
The setup conditions for efficiently sustaining nonequilibrium chemical reaction fluxes can be directly determined from these characteristics.
This is a second-order reaction. An apparent first-order constant (that depends on the global reactant concentrations) can be evaluated from $k_{0}^{+}$ to $k_{0,\text{app}}=1.4\times 10^{-6}$ s-1 assuming a concentration $c_{0}=10^{-3}$ M.
The diffusion coefficients of short oligopeptides[40, 41] are approximately equal to $D_{0}=5\times 10^{-10}$ m2.s-1, and will be assumed that they are identical (that is, $\lambda=1$).
Under such conditions, the critical parameters of this reaction-diffusion system are $\mathcal{L}_{0}=1.3$ cm and $\mathcal{T}_{0}=17$C. In order to optimize this chemical system for efficient reaction-diffusion cycles, an optimal setup should be on the centimeter scale, with a mean temperature of $40$C. Nonlinear behavior should be obtained when subjected to a gradient of at least ±$\pm 17$C.
@@
  • $l_{0}=1$ cm: $w_{\text{cold}}=3.8$, $w_{\text{warm}}=0.57$. Both lengths are larger than the length of each cold/warm region; this corresponds to a small system in which the activity is fully extended to the entire system, dominated by dissipative reactions and small concentration gradients.
  • $l_{2}=10$ cm: $w_{\text{cold}}=0.38$, $w_{\text{warm}}=0.057$. Both lengths are smaller than the length of each cold/warm region; this corresponds to a large system where the activity is restricted in the region close to the $\bar{x}=0.5$ boundary, dominated by dissipative transport and large concentration gradients.

### v3

@@ II.2.5 Optimal conditions
Overall, this implies that each process should be fast and symmetrical at mean temperature $T_{0}$. The corresponding simultaneous optimization of $\alpha$ and $\rho$ implies the maximization of the entropy production $\sigma$ (Eq. (4ap)), which represents the part of entropy production generated by the targeted chemical reaction-diffusion process. In practice, this represents only a fraction of the thermal entropy production (see Appendix B).
Moreover, it is possible to focus on a specific process to be optimized, i.e. either the reaction or the transport process, via the $\sigma_{r}$ or the $\sigma_{d}$ values. The optimization of the reaction process via $\sigma_{r}$ (see Eqs. (4bvbwcw) and Fig. S1 in the SI) implies that:
@@
A similar reasoning can be used to optimize the transport process via $\sigma_{d}$ (see Eqs. (4bvbwcx)):
@@ II.3.1 Model description
The spatial extension of the system was considered by introducing the chemical diffusion of all compounds. We limited our study to closed one-dimensional systems. A stationary temperature gradient $\theta(\bar{x})$ was imposed over the entire system. This corresponds to an unstirred system, with free internal diffusion of chemical compounds, without matter exchange with the surroundings. The system was described using a set of partial differential equations (see Eqs. (4bvbwcy)), which was numerically solved for a wide range of parameters, with a focus on determining its steady state (see Appendix A). The characteristic parameter $l_{0}$ must be introduced to account for the spatial extension, describing a one dimensional system:
$\bar{x}=\frac{x}{l_{0}}\in[0,1]$ (4au)
Chemical diffusion is described based on the diffusion parameter $d_{i}$ for each compound $U_{i}$, which is nondimensionalized by the characteristic parameter $D_{0}$ corresponding to the average diffusion constant of the chemical reactants. Consistently with the description of the chemical exchange kinetics of the previous model, chemical diffusion is described as a function of the geometric mean of the diffusion constant $d$ and the exchange ratio $\lambda$:
$d_{1}$ $=\lambda d$ (4ava)
$d_{2}$ $=\lambda^{-1}d$ (4avb)
In each point of the system, the reaction processes is still characterized as (see Eq. (4bvbwcza)):
$\sigma_{r}$ $=\varphi_{r}\cdot\alpha_{r}$ (4aw)
where the chemical flux $\varphi_{r}$ and force $\alpha_{r}$ are defined as in the two compartment system (Eqs. (4bvbwcba)-(4bvbwcbb)), on each point of the system.
The transport process of compound $i$ is now expressed as [36, 37] (see Eq. (4bvbwcza)):
$\sigma_{d}$ $=j_{i}\cdot\alpha_{d}$ (4axa)
$\textrm{with}\qquad j_{i}$ $=d_{i}\left(\frac{\partial u_{i}}{\partial\bar{x}}\right)$ (4axb)
$\textrm{and}\qquad\alpha_{d,i}$ $=\frac{1}{u_{i}}\left(\frac{\partial u_{i}}{\partial\bar{x}}\right)$ (4axc)
$j_{i}$ is the transport flux of compound $i$, and $\alpha_{d,i}$ the corresponding transport force.
The characteristic time $t_{0}$ is additionally fixed by $l_{0}$ and $D_{0}$ (see Table. 1). All parameters corresponding to chemical reactions are defined as in the stirred compartment model; however, they are now functions of position $\bar{x}$.
@@
$w$ $=\sqrt{\frac{\lambda^{\prime}d}{2\rho}}$ (4ay)
@@
  • The Soret effect was neglected, which is expected to be satisfactory for systems based on small molecular reactants [38]. However, emergent thermophoretic behaviors that originate from the coupling of chemical reactions and temperature gradients [36, 34] may occur within the framework of this model, and were specifically studied.
@@
Figure 4: Steady state spatial profiles of chemical flux of reaction $\varphi$ (red) and entropy production (solid purple, $\sigma_{r}$ and dotted purple, $\sigma_{d}$) as a function of the reaction-diffusion length $w$ as defined in Eq. (4ay), and the characteristic chemical force $\alpha_{0}$ as defined in Eq. (4bba). Model parameters: $\kappa=1$, $\lambda=1$, and $\gamma=0.2$; $d=1$, $\rho=10^{-2}$ ($w=7$, fast diffusion: A, E), $d=1$, $\rho=10^{2}$ ($w=0.07$, intermediate: B, D, F, H) and $d=10^{-2}$, $\rho=10^{2}$ ($w=0.007$, fast reaction: C, G); $\beta=1$, $\theta_{g}=10^{-1}$ ($\alpha_{0}=0.04$, linear regime: A-C, E-G) and $\beta=4$, $\theta_{g}=0.5$ ($\alpha_{0}=0.8$, far from equilibrium: D, H).
@@ II.3.2 Steady state flux profiles
When $w>0.5$, the reaction-diffusion length is larger than the sizes of the cold and warm regions $\bar{x}\in[0,0.5]$ and $\bar{x}\in[0.5,1]$. Both are homogeneous; the full system is then equivalent to the stirred-compartment model. This system is characterized by fast diffusion throughout the system, whereas the chemical reaction flux and dissipation are evenly distributed throughout the entire system (see Fig. 4A).
When $w<0.5$, an exponential decrease is observed in the values of $\varphi$ and $\alpha$, from their boundary values at $\bar{x}=0.5$ to $x=0$ or $x=1$, with a half-length equal to $w$. If $w\ll 1$, the activity of the system is restricted to a short region close to the boundary, most of the system being in equilibrium and characterized by $\varphi\approx 0$, $\alpha\approx 0$, $\sigma\approx 0$ (see Fig. 4C).
In both cases, a reaction-diffusion cycle is established, with chemical reaction and diffusion occurring simultaneously in the same spatial region, at the system boundary (see Fig. 3A, more details are available in Fig. S4).
The spatial extension of chemical activity in each region becomes asymmetric when the intensity of the temperature gradient is sufficiently high. This can be described by considering the difference in the kinetic rate in each region based on Eq. (4ada)-(4adb), as:
$\rho_{\textrm{cold}}$ $=\rho e^{-\beta\frac{\theta_{g}}{1-\theta_{g}}}$ $\text{for}~{}\bar{x}\in[0,0.5]$ (4az)
$\rho_{\textrm{warm}}$ $=\rho e^{\beta\frac{\theta_{g}}{1+\theta_{g}}}$ $\text{for}~{}\bar{x}\in[0.5,1]$ (4ba)
@@
When $\beta\theta_{g}>1$, the system becomes highly nonlinear, with a large behavioral asymmetry between each region. Typically, situations with $w_{\text{cold}}>1$ and $w_{\text{warm}}\ll 1$ can be obtained, with a homogeneous system in the cold zone and restricted to a short region in the warm zone (see Fig. 4D).
@@
  • Fast-diffusion systems (i.e. $w\gg 1$, see Fig. 4E) are characterized by a linear chemical flux profile. The system is more active at each extremity of the system, and cancels close to $\bar{x}=\nicefrac{{1}}{{2}}$, and the entropy production is dominated by the chemical reaction.
  • In fast reaction systems (i.e. $w\ll 1$, see Fig. 4G), the central inactive region becomes much larger. The chemical reaction flux and entropy production by the chemical reaction is restricted at each extremity of the system, in a region of length $w_{\text{cold}}$ at $\bar{x}\approx 0$ and $w_{\text{warm}}$ at $\bar{x}\approx 1$. In turn, the entropy production is dominated by the diffusion, and the chemical transport spreads, linking these two extremities see Fig. 3C.
This general behavior implies a spatial separation de facto of the different chemical processes. Three regions can be identified: cold and warm regions at each extremity, of half-lengths equal to $w_{\text{cold}}$ and $w_{\text{warm}}$, separated by a central region of length approximately equal to $1-2w$. The cold and warm regions are dominated by chemical reactions ($\sigma\approx\sigma_{r}$), and the central region is dominated by chemical transport ($\sigma\approx\sigma_{d}$). Thus, a spatially extended reaction-diffusion cycle is set in motion; the compounds are transported across the entire system between each chemically active region, the central transport region shrinking with $w$ (see Fig. 4E-G, and Fig. S5 for more details). This is in sharp contrast with the Heaviside gradient, where both chemical reactions and transport are effective at the same spatial locations, close to the temperature boundary (see Fig. 4A-C).
Finally, the high asymmetrization of the profile induced by high $\beta\theta_{g}$ values leads to nontrivial and complex flux profiles (see Fig. 4H). The central region around $\bar{x}=\nicefrac{{1}}{{2}}$ is characterized by a non-negligible positive flux, corresponding to a low-activity extension of the warm region. Consequently, the spatial position of the zero flux is localized at $\bar{x}<\nicefrac{{1}}{{2}}$; the cold region of the negative flux shrinks and the warm region of the positive flux stretches.
@@
The behaviors of both the linear and Heaviside profiles can be recognized (see Fig. 3B, and Fig. S6 for more details). The reaction-diffusion cycles are now in motion, with two chemical reaction zones located at $\bar{x}_{1}$ and $\bar{x}_{2}$, connected by an active chemical transport region across $\bar{x}\in[\bar{x}_{1},\bar{x}_{2}]$.
It appears that controlling the precise shape of an imposed temperature gradient enables precise control of the shape of the induced reaction-diffusion cycle. Steady dissipative chemical reactions can be localized in regions characterized by temperature inflection points (i.e., high values of $\nicefrac{{\partial^{2}T}}{{\partial x^{2}}}$), whereas chemical transport can be sustained between these active regions along temperature gradients (i.e., following high values of $\nicefrac{{\partial T}}{{\partial x}}$).
II.3.3 Steady state characteristics
@@
The values of the steady chemical flux $\varphi$, chemical force $\alpha$ and entropy production $\sigma$ can be expressed as functions of the following characteristic values:
$\alpha_{0}$ $=2\beta\gamma\theta_{g}$ (4bba)
$\varphi_{0}$ $=\kappa^{\prime}\beta\gamma\theta_{g}\rho$ (4bbb)
$\sigma_{0}$ $=2\kappa^{\prime}\beta^{2}\gamma^{2}\theta_{g}^{2}\rho$ (4bbc)
@@
Table 2: First order values for $\varphi$, $\alpha$ and $\sigma$ for low values of $\beta\theta$, as identified from numerical models (see Fig. S4-S6 in SI). The max index correspond to maximal values located at the cold/warm interface at $\bar{x}=\nicefrac{{1}}{{2}}$ for the Heaviside profile, at the system extremities $\bar{x}=0$ and $\bar{x}=1$ for the linear profile, and the inflection point of the temperature profile at $\bar{x}=\nicefrac{{1}}{{4}}$ and $\bar{x}=\nicefrac{{3}}{{4}}$. The integral values correspond to the integration of the corresponding variable on the full system $\bar{x}\in[0,1]$ for the $\sigma$ variables, or on the region restricted to the positive values of $\varphi$ and $\alpha$ (i.e. $\bar{x}\in[\bar{x}_{0},1]$, where $\varphi>0$ or $\alpha>0$ for $x>\bar{x}_{0}$).
@@
$\alpha_{0}$ $=\frac{2}{\kappa^{\prime}\rho}\varphi_{0}$ (4bq)
$\sigma_{0}$ $=\alpha_{0}\varphi_{0}$ (4br)
In the limit cases $w\ll 1$ and $w\gg 1$, and when the system is close to equilibrium (i.e. $\beta\theta_{g}<1$), the ratio of each variable to its reference value is only dependent on the reaction-diffusion length $w$ as defined in Eq. (4ay). The corresponding values are summarized in Table 2 (see Fig. S7 in SI for a graphical representation).
The situation $w\gg 1$ (that is, small systems) is essentially independent of scale. This implies that, as long as the dimension of the system is smaller than the length of the reaction-diffusion, the steady-state behavior is essentially described by the characteristic values of Eqs. (4bb), modulated by a constant factor that depends on the shape of the temperature gradient profile. Diffusion entropy production is the only varying parameter; it is proportional to $w^{-2}$, and thus is essentially negligible for small systems. This situation is the optimal case for extracting energy from the temperature gradient (by maximizing $\sigma_{t}$, as well as driving nonequilibrium chemical reaction (by maximizing both $\sigma_{r}$ and $y_{r}$).
@@
The intermediate case $w\sim 1$ is characterized by the maximal value of $\sigma_{r}$, that is, by optimal nonequilibrium transport. In this regime, the chemical reaction is also important, with a production of entropy equally shared in both processes for $w=1$. This situation is thus optimal for observing nonequilibrium cycles of reaction-diffusion, as both diffusion and reaction can be simultaneously efficient.
@@ II.3.4 System scaling
$\mathcal{L}_{0}$ $=\sqrt{\frac{D_{1}D_{2}}{k_{+,0}D_{2}+k_{-,0}D_{1}}}$ (4bs)
@@
$\mathcal{L}_{0}$ $=\sqrt{\frac{D_{0}}{2k_{0}}}$ (4bt)
The system spatial dimension $l_{0}$ should be adjusted accordingly. Nonequilibrium chemical reactions fluxes are obtained for $l_{0}<\mathcal{L}_{0}$, nonequilibrium chemical transport over large distances for $l_{0}>\mathcal{L}_{0}$, and nonequilibrium reaction-diffusion cycle for $l_{0}\sim\mathcal{L}_{0}$,
@@
Reversible peptide exchange reaction mediated by N-methylcysteine is a good candidate for being driven by temperature gradients [39]. It possesses high activation energies, and its equilibrium constant is $K=1$ close to the ambient temperature; its thermokinetic characteristics were experimentally obtained [40]:
$E_{a}^{+}$ $=55\times 10^{3}~{}\text{J}.\text{mol}^{-1}$ (4bva)
$E_{a}^{-}$ $=43\times 10^{3}~{}\text{J}.\text{mol}^{-1}$ (4bvb)
$\Delta_{r}H_{0}$ $=12\times 10^{3}~{}\text{J}.\text{mol}^{-1}$ (4bvc)
$k_{0}^{+}=k_{0}^{-}$ $=1.4\times 10^{-3}~{}\text{M}^{-1}.\text{s}^{-1}$ (4bvd)
$\text{at}\qquad T_{0}$ $=314~{}\text{K}$ (4bve)
The setup conditions for efficiently sustaining nonequilibrium chemical reaction fluxes for this reaction—leading to steady fluxes of polymerization-depolymerization—can be directly determined from these characteristics.
This is a second-order reaction. An apparent first-order constant (that depends on the global reactant concentrations) can be evaluated from $k_{0}^{+}$ to $k_{0,\text{app}}=1.4\times 10^{-6}$ s-1 assuming a concentration $c_{0}=10^{-3}$ M. The diffusion coefficients of short oligopeptides [41, 42] are approximately equal to $D_{0}=5\times 10^{-10}$ m2.s-1, and will be assumed that they are identical (that is,$\lambda=1$). Under such conditions, the critical parameters of this reaction-diffusion system are $\mathcal{L}_{0}=1.3$ cm and $\mathcal{T}_{0}=17$C.
@@
  • $l_{0}=1$ cm: $w_{\text{cold}}=3.8$, $w_{\text{warm}}=0.57$. Both lengths are larger than the length of each cold/warm region; this corresponds to a small system in which the activity is fully extended to the entire system, dominated by dissipative reactions and small concentration gradients. Setups below this scale thus correspond to regimes for performing simultaneously a steady polymerization process in the warmer region, and a steady depolymerization process in the colder region.
  • $l_{2}=10$ cm: $w_{\text{cold}}=0.38$, $w_{\text{warm}}=0.057$. Both lengths are smaller than the length of each cold/warm region; this corresponds to a large system where the activity is restricted in the region close to the $\bar{x}=0.5$ boundary, dominated by dissipative transport and large concentration gradients. Setups above this scale thus correspond to regimes for transporting peptides across the system, the chemical reaction being locally at equilibrium.
@@
From this analysis, the optimal setup conditions for sustaining steady polymerization/depolymerization process can thus be identified from the thermodynamic and kinetic characteristics of the system. The chemical system must be subjected to a temperature gradient centered on $T_{0}=314$ K (temperature for which $K=1$). Setting this gradient from $297$ K to $331$ K will drive the system in a nonlinear far-non-equilibrium state (as $T=T_{0}\pm\mathcal{T}_{0}$). Enforcing this gradient over a system length of $l_{0}=1$ cm (as $l_{0}<\mathcal{L}_{0}$) will then ensure to sustain continuous depolymerization-polymerization cycles.